OpenStudy (anonymous):

Solve S = 4v2 for v.(1 point)

OpenStudy (anonymous):

\[s = 4v^{2} \Rightarrow \frac{s}{4} = v^{2} \Rightarrow v = \pm \sqrt{\frac{v}{4}} = \frac{\pm \sqrt{v}}{2}\]

OpenStudy (zarkon):

\[v=\pm\frac{\sqrt{s}}{2}\]